For the distribution of charges shown below, rank
Answer #1

Va = KQ ( 1/d +1/3d -1/5d) = 17/15 KQ/d

Vb = KQ (1/d + 1/d -1 /3d) = 5/3 KQ/d

Vc = KQ (1/3d +1/d -1/d) = 1/3 KQ/d

Vd = -Va = -17/15 KQ/d

Hence

Vb>Va>Vc>Vd

For A = Left

B = Right

C = Right

D = Left